Circular No. 7497

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2000 S1 (SKIFF) Prediscovery observations made by LINEAR on Aug. 26, together with

additional astrometry on Sept. 26, are given on MPEC 2000-S60, along with the following orbital elements:

T = 2000 July 8.217 TT  ω = 305.^{o}699
e = 0.62614  Ω = 29.547 }2000.0
q = 2.48410 AU  i = 20.689
a = 6.64442 AU n^{o} = 0.057547 P = 17.13 years
1999TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Sept. 13 0^{h}40.^{m}61 −15^{o}22.^{′}3 1.588 2.550 158.^{o}1 8.^{o}5 15.1
23 0 32.32 −14 41.8 1.593 2.570 163.3  6.4 15.1
Oct. 3 0 23.71 −13 46.3 1.624 2.593 161.9  6.9 15.2
13 0 15.86 −12 35.3 1.680 2.619 155.1  9.2 15.3
23 0 09.62 −11 10.6 1.762 2.647 146.1 12.1 15.5

SUPERNOVA 2000dm IN UGC 11198 S. Jha, P. Challis, T. Matheson, and R. Kirshner, Harvard-Smithsonian

Center for Astrophysics, report that spectra of SN 2000dm (cf. IAUC 7495), obtained by Z. Balog on Sept. 25.1 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), show it to be a type-Ia supernova at maximum light. The spectra exhibit a blue continuum and Si II, S II, Fe II, Mg II, and Ca II features. Adopting the NASA/IPAC Extragalactic Database recession velocity of 4507 km/s for the host galaxy yields an expansion velocity for the supernova of 10 800 km/s, measured with the Si II (rest 635.5 nm) profile. The age of the supernova based on its spectral features (Riess et al. 1997, A.J. 114, 722) is 1 ± 2 days before maximum light. The spectrum also shows significant interstellar Na I D absorption (equivalent width = 0.1 nm) from our Galaxy, and the reddening towards this direction is estimated to be E(B −V) = 0.19 mag (Schlegel et al. 1998, Ap.J. 500, 525).

A. V. Filippenko and R. Chornock, University of California at Berkeley, also write: “Inspection of CCD spectra (range 330–1000 nm) of SN 2000dm, obtained on Sept. 26 UT with the Shane 3-m reflector at Lick Observatory, reveals that the object is a normal type-Ia supernova very close to maximum brightness. The presence of interstellar Na I absorption suggests that SN 2000dm is somewhat reddened.”

2000 September 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green